Angel Corporation uses activity-based costing to determine product costs for external financial reports. The company has provided the following data concerning its activity-based costing system:

Activity Cost Pools
(and Activity Measures)
Estimated
Overhead Cost
Machine related (machine-hours) $ 317,520
Batch setup (setups) $ 353,460
Order size (direct labor-hours) $ 250,560
Expected Activity
Activity Cost Pools Product X Product Y Total
Machine related 5,300 7,300 12,600
Batch setup 10,600 2,300 12,900
Order size 4,300 7,300 11,600

Assuming that actual activity turns out to be the same as expected activity, the total amount of overhead cost allocated to Product X would be closest to:
Answer #1

Overhead cost/activity would be=Respective cost/Respective activity

Overhead cost for:

Machine related=317520/12600=$25.2 per machine hour

Batch setups=(353460/12900)=$27.4 per setups

Order size=(250560/11600)=$21.6 per direct labor hours

Hence total overhead cost allocated to X =($25.2*5300 hours)+($27.4*10600 setups)+($21.6*4300 direct labor hours)

=$516880.